# Interpolation theorems for conjugations and applications

### v1

@@ 2.1. Main results
Since $q$ and $i$ were arbitrary, and $Wp(N_{X},N_{X}^{*})$ and $p(M_{Y},M_{Y}^{*})^{*}W$ are both conjugate-linear, we obtain $Wp(N_{X},N_{X}^{*})h=p(M_{Y},M_{Y}^{*})^{*}Wh$ for every $h\in H_{0}$. Consequently, by density of $H_{0}$ in $L_{N,X}$, we conclude that $Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W$. This completes the proof. ∎
Theorem 2.10.
Let $N\in{\mathcal{B}}(H)$ be normal, and let $X=\{x_{i}\}_{i\in I}$ and $Y=\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$. Let $M\in\{N,-N\}$ and assume that for every Borel subset $\Delta\subseteq\mathbb{C}$ and all $i,j\in I$, we have
  1. (a)
    $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle$;
  2. (b)
    $\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{M}(\Delta)x_{j},y_{i}\rangle$.
Then, there exists a conjugation $C$ on $L_{N,X\cup Y}$ such that
$CN_{X\cup Y}C=M_{X\cup Y}^{*}\quad\mbox{and}\quad Cx_{i}=y_{i}\;\;\mbox{for every }i\in I.$
Proof.
First, note that $L_{N,X\cup Y}=L_{M,X\cup Y}$; hence, $N_{X\cup Y}$ and $M_{X\cup Y}$ are both operators on $L_{N,X\cup Y}$, and so the statement $CN_{X\cup Y}C=M_{X\cup Y}^{*}$ makes sense. Additionally, without loss of generality, we may assume that $X$ and $Y$ are orthonormal subsets of $H$. In fact, by condition (a),
$r_{i}:=\|x_{i}\|=\sqrt{\langle E_{N}(\mathbb{C})x_{i},x_{i}\rangle}=\sqrt{\langle E_{M}(\mathbb{C})y_{i},y_{i}\rangle}=\|y_{i}\|$
for every $i\in I$; thus, the orthonormal subsets $\{r_{i}^{-1}x_{i}\}_{i\in I}$ and $\{r_{i}^{-1}y_{i}\}_{i\in I}$ satisfy conditions (a) and (b).
Using condition (a), we infer from Proposition 2.9 that there exists a conjugate-linear surjective isometry $W:L_{N,X}\to L_{M,Y}$ such that
(2.2) $Wx_{i}=y_{i}\;\mbox{and}\;Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W\;\mbox{ for all $i\in I$ and $p\in\mathbb{C}[x,y]$.}$
In particular,
(2.3) $WN_{X}=M_{Y}^{*}W\quad\mbox{and}\quad WN_{X}^{*}=M_{Y}W.$
Choose an arbitrary orthonormal basis $\{e_{j}\}_{j\in J}$ of $L_{N,X}$ that contains $X$ and, for each $j\in J$, put $f_{j}=We_{j}$. Then, since $W$ is a conjugate-linear surjective isometry, the set $\{f_{j}\}_{j\in J}$ must be an orthonormal basis of $L_{M,Y}$.
Note that, for each $j\in J$, there exist sequences of finite subsets $I_{j,n}\subseteq I$ and polynomials $p_{j,n,k}\in\mathbb{C}[x,y]$, for $n\in\mathbb{Z}^{+}$ and $k\in I_{j,n}$, such that
(2.4) $e_{j}=\lim_{n}\sum_{k\in I_{j,n}}p_{j,n,k}(N,N^{*})x_{k};$
hence, by (2.2),
(2.5) $f_{j}=We_{j}=\lim_{n}\sum_{k\in I_{j,n}}p_{j,n,k}(M,M^{*})^{*}y_{k}.$
Here, $\mathbb{Z}^{+}$ stands for the set of positive integers.
Claim: For all $i,j\in J$, we have $\langle e_{i},f_{j}\rangle=\langle e_{j},f_{i}\rangle$. First, note that in the light of condition (b), Lemma 2.8 gives
(2.6) $\langle p(N,N^{*})x_{i},y_{j}\rangle=\langle p(M,M^{*})x_{j},y_{i}\rangle\quad\mbox{for all $i,j\in I$ and $p\in\mathbb{C}[x,y]$}.$
Now, fix $i,j\in J$ and adopt the notations in (2.4) and (2.5); then
$\left\langle e_{i},f_{j}\right\rangle$ $=\lim_{n}\left\langle\sum_{k\in I_{i,n}}p_{i,n,k}(N,N^{*})x_{k},\sum_{l\in I_{j,n}}p_{j,n,l}(M,M^{*})^{*}y_{l}\right\rangle$
$=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{i,n,k}(N,N^{*})x_{k},p_{j,n,l}(M,M^{*})^{*}y_{l}\right\rangle$
$=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(M,M^{*})p_{i,n,k}(N,N^{*})x_{k},y_{l}\right\rangle.$
Hence, using (2.6), if $M=N$, we can see that
$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(N,N^{*})x_{l},y_{k}\right\rangle,$
and if $M=-N$, we obtain
$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(-N,-N^{*})x_{l},y_{k}\right\rangle.$
In both cases, we have
$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(M,M^{*})x_{l},y_{k}\right\rangle.$
Consequently, we obtain
$\left\langle e_{i},f_{j}\right\rangle$ $=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})x_{l},p_{i,n,k}(M,M^{*})^{*}y_{k}\right\rangle$
$=\lim_{n}\left\langle\sum_{l\in I_{j,n}}p_{j,n,l}(N,N^{*})x_{l},\sum_{k\in I_{i,n}}p_{i,n,k}(M,M^{*})^{*}y_{k}\right\rangle$
[…]
This proves the claim.
Note that $L_{N,X\cup Y}=\vee\left(L_{N,X}+L_{M,Y}\right)$; hence, $\{e_{j}\}_{j\in J}$ and $\{f_{j}\}_{j\in J}$ are both orthonormal subsets of $L_{N,X\cup Y}$. By [29, Theorem 2.1], there exists a conjugation $C$ on $L_{N,X\cup Y}$ such that $Ce_{j}=f_{j}$ for every $j\in J$. In particular, since $\{e_{j}\}_{j\in J}$ contains $X$, we get from (2.2) that
$Cx_{i}=Wx_{i}=y_{i}\quad\mbox{for every $i\in I$.}$
It remains to check that $CN_{X\cup Y}C=M_{X\cup Y}^{*}$. Noting that, for every $j\in J$, $Ce_{j}=f_{j}=We_{j}$ and $Cf_{j}=C^{2}e_{j}=e_{j}=W^{-1}f_{j}$, and taking into account that $C$, $W$ and $W^{-1}$ are conjugate-linear, it follows that $W=C_{1}$ and $W^{-1}=C_{2}$ where $C_{1}$ and $C_{2}$ are defined as
$\begin{array}[]{rcl}C_{1}:L_{N,X}&\to&L_{M,Y}\\ h&\mapsto&Ch\end{array}\quad\mbox{and}\quad\begin{array}[]{rcl}C_{2}:L_{M,Y}&\to&L_{N,X}\\ h&\mapsto&Ch\end{array}.$
Combining these facts with (2.3), we get that for every $j\in J$,
$CM_{X\cup Y}Ce_{j}=CM_{X\cup Y}We_{j}=CM_{Y}We_{j}=CWN_{X}^{*}e_{j}=C_{2}WN_{X}^{*}e_{j}=N_{X}^{*}e_{j}=N_{X\cup Y}^{*}e_{j}$
and
$CN_{X\cup Y}Cf_{j}=CN_{X\cup Y}e_{j}=CN_{X}e_{j}=C_{1}N_{X}e_{j}=WN_{X}e_{j}=M_{Y}^{*}We_{j}=M_{Y}^{*}f_{j}=M_{X\cup Y}^{*}f_{j}.$
Suppose that $M=N$; it follows that $CN_{X\cup Y}Ce_{j}=N_{X\cup Y}^{*}e_{j}$ and $CN_{X\cup Y}Cf_{j}=N_{X\cup Y}^{*}f_{j}$ for every $j\in J$. Hence, since $L_{N,X\cup Y}=\vee\left(L_{N,X}+L_{M,Y}\right)=\vee\left(\{e_{j},f_{j}\}_{j\in J}\right)$, we obtain $CN_{X\cup Y}C=N_{X\cup Y}^{*}$. On the other hand, if $M=-N$, we get $CN_{X\cup Y}Ce_{j}=-N_{X\cup Y}^{*}e_{j}$ and $CN_{X\cup Y}Cf_{j}=-N_{X\cup Y}^{*}f_{j}$ for every $j\in J$. Therefore, $CN_{X\cup Y}C=-N_{X\cup Y}^{*}$. In both cases, we have the desired result. This completes the proof of the theorem.
@@
Let $N\in\mathcal{B}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $T:H\to K$ be a bounded linear (or conjugate-linear) operator. The following statements hold:
@@
$L=\begin{bmatrix}N&0\\ 0&M\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}\quad\mbox{and}\quad A=\begin{bmatrix}0&0\\ T&0\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}.$
Then, $L$ and $A$ are bounded linear operators acting on the Hilbert space $H\oplus K$. Furthermore, one can verify that $AL=LA$. Since $L$ is normal, by the Fuglede Theorem, $AE_{L}(\Delta)=E_{L}(\Delta)A$ for every Borel subset $\Delta\subseteq\mathbb{C}$.
@@
$E_{L}(\Delta)=\begin{bmatrix}E_{N}(\Delta)&0\\ 0&E_{M}(\Delta)\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}.$
@@
Assume now that $T$ is a conjugate-linear operator such that $TN=M^{*}T$, and let $\Delta\subseteq\mathbb{C}$ be a Borel subset. Since $M$ is normal, there exists a conjugation $C$ on $K$ such that $CMC=M^{*}$. Hence, $TN=M^{*}T=(CMC)T$, and since $C^{2}=I$, we obtain
(2.7) $(CT)N=C^{2}MCT=M(CT).$
Clearly, $CT:H\to K$ is a bounded linear operator. By our previous result, we have
@@
By [8, Page 287, Lemma 8.7], $E_{M}(\Delta)=\chi_{\Delta}(M)\in W^{*}(M)$. Hence, there exists a net $\{p_{i}\}_{i}$ in $\mathbb{C}[x,y]$ such that $p_{i}(M,M^{*})$ converges to $E_{M}(\Delta)$ in the weak operator topology. One can easily check that $Cp_{i}(M,M^{*})C=p_{i}(M,M^{*})^{*}$. Consequently, taking the limit in the weak operator topology, we get $CE_{M}(\Delta)C=E_{M}(\Delta)^{*}=E_{M}(\Delta).$ Therefore, from (2.8), we obtain
@@
We are now in a position to prove Theorems 2.1 and 2.3 and 2.4.
@@
The implication (i)$\implies$(ii) follows directly from the properties of $C$ as a conjugate-linear isometry, satisfying $Cx_{i}=y_{i}$, $Cy_{i}=C^{2}x_{i}=x_{i}$ and $CT=CTC^{2}=T^{*}C$ for all $i\in I$ and $T\in\mathcal{M}$.
@@
(ii)$\implies$(iii). Let $V_{0}:{\rm ker}(V)^{\perp}\to{\rm ran}(V)$ be the invertible conjugate-linear isometry given by $V_{0}h=Vh$. Define the conjugate-linear map $V^{\#}:H\to H$ by $V^{\#}h=V_{0}^{-1}h$ if $h\in{\rm ran}(V)$ and $V^{\#}h=0$ if $h\in{\rm ran}(V)^{\perp}$. Then, $V^{\#}V$ is the orthogonal projection onto ${\rm ker}(V)^{\perp}$. Moreover, using the polarization identity, it follows that $\langle Vh,k\rangle=\langle V^{\#}k,h\rangle$ for all $h\in{\rm ker}(V)^{\perp}$ and $k\in{\rm ran}(V)$. Thus,
(2.9) $\langle Vh,k\rangle=\langle V^{\#}k,h\rangle\quad\mbox{for all }h,k\in H.$
Now, let $P\in\mathcal{M}$ be an orthogonal projection. Using the above relation, for all $i,j\in I$, we obtain
@@
(iii)$\implies$(ii). According to [10, Page 52, Lemma II.2.8], there exists a self-adjoint operator $A\in{\mathcal{B}}(H)$ such that $\mathcal{M}=W^{*}(A)$. It suffices to show that there exists a conjugation $C$ on $H$ such that $CAC=A$ and $Cx_{i}=y_{i}$ for every $i\in I$. In fact, any polynomial in $A$ is $C$-symmetric; moreover, since the class of $C$-symmetric operators is closed in the weak operator topology, every operator in $W^{*}(A)$, and thus in $\mathcal{M}$, is $C$-symmetric.
Let $\Delta\subseteq\mathbb{C}$ be a Borel subset. Then, by [8, Page 287, Lemma 8.7], the orthogonal projection $E_{A}(\Delta)$ belongs to $\mathcal{M}$. Therefore, by hypothesis, for all $i,j\in I$, we have $\langle E_{A}(\Delta)x_{i},x_{j}\rangle=\langle E_{A}(\Delta)y_{j},y_{i}\rangle$ and $\langle E_{A}(\Delta)x_{i},y_{j}\rangle=\langle E_{A}(\Delta)x_{j},y_{i}\rangle$.
For convenience, put $Z=\{x_{i},y_{i}\}_{i\in I}$, $H_{0}=L_{A,Z}$ and $H_{1}=L_{A,Z}^{\perp}$. Then, $H_{0}$ and $H_{1}$ are reducing subspaces of $A$ and $Z\subset H_{0}$. According to Theorem 2.10, there exists a conjugation $C_{0}$ on $H_{0}$ such that $C_{0}A_{|H_{0}}C=A_{0}$ and $C_{0}x_{i}=y_{i}$ for every $i\in I$. Since $A_{|H_{1}}$ is normal, there exists a conjugation $C_{1}$ on $H_{1}$ such that $C_{1}A_{|H_{1}}C_{1}=A_{|H_{1}}$. Clearly, $C=C_{0}\oplus C_{1}$ is a conjugation on $H$; additionally, $CAC=A$ and $Cx_{i}=C_{0}x_{i}=y_{i}$ for every $i\in I$. ∎
Proof of Theorem 2.3.
The implication (i)$\implies$(ii) is obvious, and the implication (ii)$\implies$(iii) follows a similar argument as (i)$\implies$(ii) in Theorem 2.1.
To show that (iii)$\implies$(iv), first note that by Lemma 2.11, for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $k\in K$, we have $VE_{N_{k}}(\Delta)=E_{N_{k}}(\Delta)V$. Hence, for any finite subset $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, we have $V\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)=\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)V$. Put $P=\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})$; the remainder of the proof follows similarly to (2.10) and (2.11).
(iv)$\implies$(i). For convenience, put $\mathcal{M}=W^{*}(N_{k}:k\in K)$. By the Fuglede Theorem, we have $N_{k}N_{l}^{*}=N_{l}^{*}N_{k}$ for any $k,l\in K$. Therefore, it is not hard to see that $\mathcal{M}$ is an abelian von Neumann algebra. Hence, by Theorem 2.1, we need to show that for all orthogonal projections $P\in\mathcal{M}$ and $i,j\in I$,
(2.12) $\langle Px_{i},x_{j}\rangle=\langle Py_{j},y_{i}\rangle\quad\mbox{and}\quad\langle Px_{i},y_{j}\rangle=\langle Px_{j},y_{i}\rangle.$
First, we show that $\mathcal{M}={\rm cl-WOT}\mathcal{M}_{0}$; i.e., the closure in the weak operator topology of
$\mathcal{M}_{0}={\rm span}\left\{\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k}):K_{0}\subseteq K\mbox{ is finite}\mbox{ and }\Delta_{k}\mbox{ is a Borel subset of }\mathbb{C}\right\}.$
By [8, Page 287, Lemma 8.7], any spectral projection $E_{N_{k}}(\Delta)=\chi_{\Delta}(N_{k})\in W^{*}(N_{k})\subseteq\mathcal{M}$. Hence, $\mathcal{M}_{0}\subseteq\mathcal{M}$, and consequently ${\rm cl-WOT}\mathcal{M}_{0}\subseteq\mathcal{M}$. Let us show the other containment. According to spectral theory, for every $k\in K$, we have
$N_{k}\in\bigvee\big{\{}E_{N_{k}}(\Delta):\Delta\mbox{ is a Borel subset of }\mathbb{C}\big{\}}.$
This implies that $N_{k}\in{\rm cl-WOT}\mathcal{M}_{0}$ for every $k\in K$. It is easy to see that ${\rm cl-WOT}\mathcal{M}_{0}$ is a self-adjoint subalgebra of ${\mathcal{B}}(H)$ that contains the identity. Moreover, since, by the Fuglede Theorem, the spectral measures $\{E_{N_{k}}\}_{k\in K}$ mutually commute and $E_{N_{k}}(\Delta)E_{N_{k}}(D)=E_{N_{k}}(\Delta\cap D)$ for any Borel subsetes $\Delta$ and $D$ of $\mathbb{C}$, one can easily show that $\mathcal{M}_{0}$, and thus ${\rm cl-WOT}\mathcal{M}_{0}$, is also multiplicative. Hence, ${\rm cl-WOT}\mathcal{M}_{0}$ is a von Neumann algebra containing $\{N_{k}\}_{k\in K}$. Therefore, by definition of $\mathcal{M}$, we must have $\mathcal{M}\subseteq{\rm cl-WOT}\mathcal{M}_{0}$.
Now, by hypothesis, any $T\in\mathcal{M}_{0}$ satisfies the equalities (2.12). Then, by taking limits, in the weak operator topology, of operators in $\mathcal{M}_{0}$, we obtain that (2.12) is true for every $T\in\mathcal{M}$; in particular, for orthogonal projections in $\mathcal{M}$. ∎
@@
The implication (iv)$\implies$(i) follows from Theorem 2.3 (iv)$\implies$(i). In fact, for any finite subset $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, the orthogonal projections $\{E_{N_{k}}(\Delta_{k})\}_{k\in K_{0}}$ are mutually commuting. Thus, their product $P$ is also an orthogonal projection; which gives
@@
(i)$\implies$(v). Clearly, $\|x\|=\|Cx\|=\|y\|$. It is easy to check that $C(x\otimes y)C=(Cx)\otimes(Cy)$; hence, for any $k\in K$ and $\lambda\in\mathbb{C}$, we have

### v3

@@ 2.1. Main results
Since $n$, $q$, $\{k_{l}\}_{1\leq l\leq n}$ and $i$ were arbitrary, and taking into account that $T$ and $S$ are both conjugate-linear, we obtain $Th=Sh$ for every $h\in\mathcal{H}_{0}$. Consequently, the desired equality follows by density of $\mathcal{H}_{0}$ in $L_{\mathcal{A},A}$. ∎
@@
Let $N\in\mathcal{B}(\mathcal{H})$ and $M\in\mathcal{B}(\mathcal{K})$ be normal operators, and let $T:\mathcal{H}\to\mathcal{K}$ be a bounded linear (or conjugate-linear) operator. The following statements hold:
@@
$L=\begin{bmatrix}N&0\\ 0&M\end{bmatrix}\begin{array}[]{l}\mathcal{H}\\ \mathcal{K}\end{array}\quad\mbox{and}\quad A=\begin{bmatrix}0&0\\ T&0\end{bmatrix}\begin{array}[]{l}\mathcal{H}\\ \mathcal{K}\end{array}.$
Then, $L$ and $A$ are bounded linear operators acting on the Hilbert space $\mathcal{H}\oplus\mathcal{K}$. Furthermore, one can verify that $AL=LA$. Since $L$ is normal, by the Fuglede Theorem, $AE_{L}(\Delta)=E_{L}(\Delta)A$ for every Borel subset $\Delta\subseteq\mathbb{C}$.
@@
$E_{L}(\Delta)=\begin{bmatrix}E_{N}(\Delta)&0\\ 0&E_{M}(\Delta)\end{bmatrix}\begin{array}[]{l}\mathcal{H}\\ \mathcal{K}\end{array}.$
@@
Assume now that $T$ is a conjugate-linear operator such that $TN=M^{*}T$, and let $\Delta\subseteq\mathbb{C}$ be a Borel subset. Since $M$ is normal, there exists a conjugation $C$ on $\mathcal{K}$ such that $CMC=M^{*}$. Hence, $TN=M^{*}T=(CMC)T$, and since $C^{-1}=C$, we obtain
(2.3) $(CT)N=M(CT).$
Clearly, $CT:\mathcal{H}\to\mathcal{K}$ is a bounded linear operator. By our previous result, we have
@@
By [9, p. 287, Lemma 8.7], $E_{M}(\Delta)=\chi_{\Delta}(M)\in W^{*}(M)$, the von Neumann algebra generated by $M$. Hence, there exists a net $\{p_{i}\}_{i}$ in $\mathcal{P}_{2}(\mathbb{C})$ such that $p_{i}(M,M^{*})$ converges to $E_{M}(\Delta)$ in the weak operator topology. One can easily check that $Cp_{i}(M,M^{*})C=p_{i}(M,M^{*})^{*}$ for any $i$. Consequently, taking the limit in the weak operator topology, we get $CE_{M}(\Delta)C=E_{M}(\Delta)^{*}=E_{M}(\Delta).$ Therefore, from (2.4), we obtain
@@
We are now in a position to prove Theorems 2.1 and 2.4 and 2.8.
@@
(i)$\implies$(ii). Adopt the notations in Theorem 2.1 (i). Since, for every $k\in K$, $CN_{k}=M_{k}^{*}C$ and $CM_{k}=N_{k}^{*}C$, it follows by Lemma 2.15 that for all $k\in K_{0}$ and $l\in K_{1}$, $CE_{M_{k}}(\Delta_{k})=E_{N_{k}}(\Delta_{k})C$ and $CE_{N_{l}}(D_{l})=E_{M_{l}}(D_{l})C$. Hence,
$\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)C=C\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)\quad\mbox{and}\quad\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)C=C\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right).$
Thus, taking into account the equalities $Cx_{i}=y_{i}$ and $Cy_{i}=x_{i}$, we get
$\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)y_{j}\right\rangle$ $=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)Cy_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)Cx_{j}\right\rangle$
$=\left\langle C\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)y_{i},C\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{j}\right\rangle$
$=\left\langle\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{j},\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)y_{i}\right\rangle.$
The other equality can be obtained in a similar way.
(ii)$\implies$(i). Put $K_{1}=K_{2}=K$ and let $\widetilde{K}$ be the disjoint union of $K_{1}$ and $K_{2}$. Let $\mathcal{A}=\{A_{k}\}_{k\in\widetilde{K}}$ and =$\mathcal{B}=\{B_{k}\}_{k\in\widetilde{K}}$ where
$A_{k}=\begin{cases}N_{k}&\text{ if }k\in K_{1};\\ M_{k}&\text{ if }k\in K_{2},\end{cases}\quad\mbox{and}\quad B_{k}=\begin{cases}M_{k}&\text{ if }k\in K_{1};\\ N_{k}&\text{ if }k\in K_{2}.\end{cases}$
Now, define $\widetilde{I}$ to be the disjoint union of $I_{1}$ and $I_{2}$ where $I_{1}=I_{2}=I$, and let $A=\{a_{i}\}_{i\in\widetilde{I}}$ and $B=\{b_{i}\}_{i\in\widetilde{I}}$ be such that
$a_{i}=\begin{cases}x_{i}&\text{ if }i\in I_{1};\\ y_{i}&\text{ if }i\in I_{2},\end{cases}\quad\mbox{and}\quad b_{i}=\begin{cases}y_{i}&\text{ if }i\in I_{1};\\ x_{i}&\text{ if }i\in I_{2}.\end{cases}$
Since, for all subsets $\{k_{l}\}_{l=1}^{4}\subseteq K$ and Borel subsets $\{\Delta_{k_{l}}\}_{l=1}^{4}$ of $\mathbb{C}$, the orthogonal projections $E_{N_{k}}(\Delta_{k_{1}})$, $E_{N_{k}}(\Delta_{k_{2}})$, $E_{M_{k}}(\Delta_{k_{3}})$ and $E_{M_{k}}(\Delta_{k_{4}})$ are mutually commuting, one can easily see that $\mathcal{A}=\{A_{k}\}_{k\in\widetilde{K}}$, =$\mathcal{B}=\{B_{k}\}_{k\in\widetilde{K}}$, $A$ and $B$ satisfy the conditions in Proposition 2.14. Consequently, and since $L_{\mathcal{A},A}=L_{\mathcal{B},B}=L_{\mathcal{N}\cup\mathcal{M},X\cup Y}=\mathcal{H}$, there exists a conjugate-linear surjective isometry $C:\mathcal{H}\to\mathcal{H}$ such that
(2.5) $Ca_{i}=b_{i}\quad\mbox{and}\quad Cp(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})=p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}C$
for all $i\in I$, $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{1\leq l\leq n}\subseteq\widetilde{K}$. Furthermore, by definition of $\mathcal{A}=\{A_{k}\}_{k\in\widetilde{K}}$, =$\mathcal{B}=\{B_{k}\}_{k\in\widetilde{K}}$, $A$ and $B$, the roles of $A_{k_{l}}$ (resp. $a_{i}$) and $B_{k_{l}}$ (resp. $b_{i}$) can be interchanged in (2.5); that is,
(2.6) $Cb_{i}=a_{i}\quad\mbox{and}\quad Cp(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})=p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})^{*}C.$
In particular, we have
$Cx_{i}=y_{i}\quad\mbox{and}\quad CN_{k}=M_{k}^{*}C\quad\mbox{for every $k\in K$}.$
Now, we only need to show that $C$ is a conjugation. Let us first show that $C^{2}=I_{\mathcal{H}}$. Since $\mathcal{H}=L_{\mathcal{A},A}$, it suffices to show that $C^{2}h=h$ for every vector $h$ of the form $h=p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}$ where $i\in I$, $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{1\leq l\leq n}\subseteq\widetilde{K}$. Let $h$ be such vector; then, (2.5) and (2.6) imply
$C^{2}h=C^{2}p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}$ $=Cp(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}b_{i}$
$=p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}=h.$
Finally, the isometric property in Definition 1.1 (ii) can be obtain from (2.7) since $C^{\#}=C^{-1}=C$. ∎
Proof of Theorem 2.4.
The implication (i)$\implies$(ii) is obvious.
@@
(ii)$\implies$(iii). Let $V_{0}:{\rm ker}(V)^{\perp}\to{\rm ran}(V)$ be the invertible conjugate-linear isometry given by $V_{0}h=Vh$. Define the conjugate-linear map $V^{\#}:\mathcal{H}\to\mathcal{H}$ by $V^{\#}h=V_{0}^{-1}h$ if $h\in{\rm ran}(V)$ and $V^{\#}h=0$ if $h\in{\rm ran}(V)^{\perp}$. Then, $V^{\#}V$ is the orthogonal projection onto ${\rm ker}(V)^{\perp}$. Moreover, using the polarization identity, it follows that $\langle Vh,k\rangle=\langle V^{\#}k,h\rangle$ for all $h\in{\rm ker}(V)^{\perp}$ and $k\in{\rm ran}(V)$. Thus,
(2.7) $\langle Vh,k\rangle=\langle V^{\#}k,h\rangle\quad\mbox{for all }h,k\in\mathcal{H}.$
Let $K_{0}\subseteq K$ be a finite subset, and let $\{\Delta_{k}\}_{k\in K_{0}}$ be Borel subsets of $\mathbb{C}$. In view of Lemma 2.15, for all $k\in K_{0}$, we have $VE_{N_{k}}(\Delta_{k})=E_{N_{k}}(\Delta_{k})V$. Hence, $V\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)=\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)V$. Put $P=\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})$. Thus,
@@
(iii)$\implies$(i). Assume that $M_{k}=N_{k}$ for every $k\in K$. For convenience, put $\mathcal{L}_{0}=L_{\mathcal{N},X\cup Y}$ and $\mathcal{L}_{1}=\mathcal{L}_{0}^{\perp}$. Then, $\mathcal{L}_{0}$ reduces $N_{k}$ for every $k\in K$; additionally, $x_{i},y_{i}\in\mathcal{L}_{0}$ for all $i\in I$. It is not hard to see that $\left\{{N_{k}}_{|\mathcal{L}_{0}}\right\}_{k\in K}$, $\left\{{M_{k}}_{|\mathcal{L}_{0}}\right\}_{k\in K}$, $X$ and $Y$ satisfy assertion (ii) of Theorem 2.1. Thus, by the same theorem, there exists a conjugation $C_{0}$ on $\mathcal{L}_{0}$ such that $C_{0}{N_{k}}_{|\mathcal{L}_{0}}C_{0}=\left({N_{k}}_{|\mathcal{L}_{0}}\right)^{*}$ and $C_{0}x_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
@@
On the other hand, since the normal operators ${N_{k}}_{|\mathcal{L}_{1}}$ are mutually commuting, they generate an abelian von Neumann algebra in $\mathcal{B}(\mathcal{L}_{1})$. Hence, according to [11, p. 52, Lemma II.2.8], there exists a self-adjoint operator $A\in\mathcal{B}(\mathcal{L}_{1})$ such that ${N_{k}}_{|\mathcal{L}_{1}}\in W^{*}(A)$ for every $k\in K$. By self-adjointness of $A$, there exists a conjugation $C_{1}$ on $\mathcal{L}_{1}$ such that $C_{1}AC_{1}=A$. Thus, any polynomial in $A$ is $C_{1}$-symmetric; moreover, since the class of $C_{1}$-symmetric operators is closed in the weak operator topology, we get
(2.10) $C_{1}{N_{k}}_{|\mathcal{L}_{1}}C_{1}=\left({N_{k}}_{|\mathcal{L}_{1}}\right)^{*}\quad\mbox{for every $k\in K$}.$
Therefore, one can readily see that $C=C_{0}\oplus C_{1}$ is a conjugation on $\mathcal{H}=\mathcal{L}_{0}\oplus\mathcal{L}_{1}$ that satisfies all desired equalities. ∎
@@
The implication (iv)$\implies$(i) follows from Theorem 2.4 (iii)$\implies$(i). In fact, for any finite subset $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, the orthogonal projections $\{E_{N_{k}}(\Delta_{k})\}_{k\in K_{0}}$ are mutually commuting. Thus, their product $P$ is also an orthogonal projection; which gives
(i)$\implies$(v). In view of (1.1), for any $k\in K$ and $\lambda\in\mathbb{C}$, we have